OpenStudy (chaise):

Could anyone try and help me with a question? sqrt(8+sqrt(8+sqrt(8+sqrt(8. All the way to inifity?

OpenStudy (chaise):

\[\sqrt{8+\sqrt{8+\sqrt{8+\sqrt{8+\sqrt{8+.....}}}}}\] I know you have to make it equal to x, and then solve somehow.. but I can't quite remember, it's been about a year since doing this.

OpenStudy (chaise):

I know the answer, I want to know HOW. Thanks in advance :)

OpenStudy (anonymous):

x^2=8+x

OpenStudy (chaise):

This gives: x^2-x-8=0 -b+sqrt(b^2-4ac)/2a to solve? Looks correct, but how did you get there? Steps please? If you can.

OpenStudy (chaise):

I got it, thanks anyway. :) Appreciated. Just needed a hint